Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
  Machine-hours 26,000 36,000     62,000
  Fixed manufacturing overhead costs $ 730,000 $ 300,000     $ 1,030,000
  Variable manufacturing overhead per machine-hour $ 5.70 $ 5.70    

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

  Job D-70: Molding Fabrication Total
  Direct materials cost $ 378,000 $ 326,000 $ 704,000
  Direct labor cost $ 210,000 $ 120,000 $ 330,000
  Machine-hours 18,000 8,000 26,000

  

  Job C-200: Molding Fabrication Total
  Direct materials cost $ 230,000 $ 300,000 $ 530,000
  Direct labor cost $ 130,000 $ 250,000 $ 380,000
  Machine-hours 8,000 28,000 36,000

  
Delph had no overapplied or underapplied manufacturing overhead during the year.

Assume Delph uses a plantwide overhead rate based on machine-hours.

  

Required:
1-a.

Compute the predetermined plantwide overhead rate. (Round your answer to 2 decimal places.)

          

1-b.

Compute the total manufacturing costs assigned to Job D-70 and Job C-200. (Round your intermediate calculations to 2 decimal places.)

          

1-c.

If Delph establishes bid prices that are 140% of total manufacturing costs, what bid price would it have established for Job D-70 and Job C-200? (Round your intermediate calculations to 2 decimal places.)

          

1-d.

What is Delph’s cost of goods sold for the year? (Round your intermediate calculations to 2 decimal places.)

       

Assume Delph uses departmental overhead rates based on machine-hours.

  

2-a.

Compute the predetermined departmental overhead rates. (Round your answer to 2 decimal places.)

         

2-b.

Compute the total manufacturing costs assigned to Job D-70 and Job C-200. (Round your intermediate calculations to 2 decimal places.)

          

2-c.

If Delph establishes bid prices that are 140% of total manufacturing costs, what bid price would it have established for Job D-70 and Job C-200? (Round your intermediate calculations to 2 decimal places.)

          

2-d.

What is Delph’s cost of goods sold for the year? (Round your intermediate calculations to 2 decimal places.)

       

Please show how you got the answers
Answer #1

1 (a )-

statement of total overhead-

particulars molding fabrication total
variable rate 5.7 5.7
no. of mach. hrs. 26000 36000 62000
variable overhead 148200 205200 353400
fixed overhead 730000 300000 1030000
total overheads 878200 505200 1383400

overhead rate= total manufacturing overhead/no. of machine hrs.

= 1383400/62000 = 22.31

1(b)-

m. overhead assigned to job D70 = 1383400/62000)*26000 = 580135.48

m. overhead assigned to job C-200 = 1383400/62000)*36000 = 803264.52

total cost for D70 = 704000+330000+580135.48 = 1614135.48

total cost to C200 = 530000+380000+803264.52 = 1713264.52

1(c)-

bid price for job D70 = 140%*1614135.48 = 2259789.67

bid price for job C200 = 140%*1713264.52 = 2398570.33

1(d)-

cost of goods sold = begining inventory+ direct material+ direct labour+ overhead - closing inventory

= 0 +(704000+530000)+(330000+380000)+1383400-0 = 3327400

2(a)-

statement of total overhead-

particulars molding fabrication total
variable rate 5.7 5.7
no. of mach. hrs. 26000 36000 62000
variable overhead 148200 205200 353400
fixed overhead 730000 300000 1030000
total overheads 878200 505200 1383400

overhead rate department wise= total manufacturing cost/no. of machine hrs.

molding = 878200/26000 = 33.78

fabrication = 505200/36000 = 14.03

2(b) -

total cost to job D70 = 18000* 33.78 + 8000*14.03 + 878200 = 1598480

total cost to job C200 = 8000*33.78 + 28000*14.03 + 505200 = 1168280

2(c) -

bid price for job D70 = 140%*1598480 = 2237872

bid price for job C200 = 140%*1168280 =1635592

2(d) -

cost of goods sold = begining inventory+ direct material+ direct labour+ overhead - closing inventory

  = 0 +(704000+530000)+(330000+380000)+1383400-0 = 3327400